OpenStudy (erinkb99):

SIMPLE DIFFERENTIAL EQUATION: Solve the differential equation dy/dx= y^2/x^3 for y = f(x) with the condition y(1) = 1.

OpenStudy (phi):

do you know how to "separate the variables" i.e. move all the x's (and dx's) on one side and the y;s on the other ?

OpenStudy (erinkb99):

yes. I did it, does this look ok?

OpenStudy (erinkb99):

dy/dx=y2(1/x3) 1/y2 dy= 1/x3 dx @phi

OpenStudy (phi):

ok, and if you write it with negative exponents \[ y^{-2} \ dy = x^{-3} \ dx \] now integrate both sides \[ \int y^{-2} \ dy = \int x^{-3} \ dx \]

OpenStudy (erinkb99):

-y^-1 = -1/2x^-2 + C ? @phi

OpenStudy (phi):

yes. you can probably clean it up a bit

OpenStudy (phi):

remember y(1)= 1 that means x=1, y=1 is a solution. So you can solve for C

OpenStudy (erinkb99):

I'm having a hard time getting the y truly by itself. Can you show me that?

OpenStudy (phi):

I would not bother. we know the point (1,1) is a solution (that is what f(1)= 1 means) so replace x with 1 and y with 1 solve for C

OpenStudy (erinkb99):

-1^-1 = -1/2(1)^-2 + C -1= -0.5 + C -0.5=C

OpenStudy (phi):

ok. \[ - \frac{1}{y}= - \frac{1}{2x}- \frac{1}{2} \] I would multiply both sides by -1 to get rid of all the minus signs \[ \frac{1}{y}= \frac{1}{2x}+ \frac{1}{2} \] Next, add the two fractions on the right side

OpenStudy (phi):

that should be x^2 (not x) in the bottom

OpenStudy (phi):

the last step would be to "flip" i.e. invert both sides to get y = some fraction of x

OpenStudy (erinkb99):

so what does C equal?

OpenStudy (phi):

C is the unknown constant of integration. You found it equal to -0.5 i.e. - ½ we put in -½ in to your solution now we simplify /rewrite to make it look a bit nicer.

OpenStudy (erinkb99):

Ok great! Thanks so much. I have some more questions I would like you to check. I'll post them separately. I'll fan you.

OpenStudy (phi):

in other words, your solution is \[ - \frac{1}{y}= - \frac{1}{2x^2}- \frac{1}{2} \\ \frac{1}{y}= \frac{1}{2x^2} +\frac{x^2}{2x^2}\\ y = \frac{2x^2}{x^2+1} \]

OpenStudy (erinkb99):

Yep!

OpenStudy (erinkb99):

Thanks again.